### v1

@@ 5 A thick-restart IPJDSVD algorithm with deflation and purgation
For practical purpose, we must limit the subspace dimension $k\leq k_{\max}$. If JDSVD and IPJDSVD do not yet converge until $k=k_{\max}$, we must restart them. Furthermore, if we are required to compute $\ell>1$ singular triplets of $A$ corresponding to the $\ell$ singular values $\sigma_{1},\ldots,\sigma_{\ell}$ closest to the target $\tau$ (cf. (1.2)), an appropriate deflation is necessary. In addition, a certain purgation is crucial and may greatly reduce the outer iterations for computing the second to the $\ell$th singular triplets.
@@ 5.1 A new thick-restart
If IPJDSVD does not yet converge for computing the desired $\ell$ singular triplets for a given $k_{\max}$, we modify a commonly used thick restart, initially proposed in [25] and later popularized in the JDSVD and JDGSVD type methods [10, 11, 12, 29]. In what follows we propose a more effective thick-restart IPJDSVD than the standard thick-restart JDSVD. Our thick-restart retains some $k_{\mathrm{new}}$-dimensional restarting subspaces with
(5.1) $k_{\mathrm{new}}=\max\{k_{\min},\tilde{m}\}$
where $k_{\min}>1$ is a user-prescribed minimal dimension of the searching subspaces and $\tilde{m}$ is the number of approximate singular triplets of $A$ that are selected to participate in the preconditioned correction equation (4.10) in the current expansion phase.
Specifically, assume that $\tilde{m}$ approximate singular triplets $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i}),i=1,i_{2},\dots$, $i_{\tilde{m}}$ of $A$ with $i_{2},\dots,i_{\tilde{m}}\in\{2,\dots,k\}$ join the preconditioned (4.10) before restart. If $\tilde{m}\leq k_{\min}$, we simply restore the standard thick-restart scheme, which forms the orthonormal basis matrices of the new $k_{\min}$ dimensional $\mathcal{U}$ and $\mathcal{V}$ by
(5.2) $\widetilde{U}=[\tilde{u}_{1},\tilde{u}_{2},\dots,\tilde{u}_{k_{\min}}],\qquad\widetilde{V}=[\tilde{v}_{1},\tilde{v}_{2},\dots,\tilde{v}_{k_{\min}}],$
@@
[…]
If $\tilde{m}>k_{\min}$, we form the orthonormal basis matrices of the new $\tilde{m}$ dimensional $\mathcal{U}$ and $\mathcal{V}$ and the corresponding projection matrix of $A$ by
[…]
with $\widetilde{U}_{\tilde{m}}$ and $\widetilde{V}_{\tilde{m}}$ defined by (4.40). In this way, the $k_{\rm new}$-dimensional restarting subspaces contain approximate left and right singular subspaces associated with all available approximate singular values clustered at $\tau$. Whenever $m<k_{\max}$ considerably, the new thick-restart IPJDSVD is expected to substantially outperform the standard thick-restart JDSVD in terms of the outer iterations.
@@ 5.2 Deflation and purgation
Suppose that we are required to compute the $\ell>1$ singular triplets of $A$ corresponding to the $\ell$ singular values closest to the target $\tau$. Then IPJDSVD combined with the deflation advocated in [10] meet this demand perfectly. Concretely, assume that $j<\ell$ approximate singular triplets $\left(\theta_{i,\mathrm{c}},u_{i,\mathrm{c}},v_{i,\mathrm{c}}\right)$ have already converged to $(\sigma_{i},u_{i},v_{i})$, $i=1,\dots,j$ with the residual $r^{(i,\mathrm{c})}=r(\theta_{i,\mathrm{c}},u_{i,\mathrm{c}},v_{i,\mathrm{c}})$ defined by (2.4) satisfying
(5.5) $\|r^{(i,\mathrm{c})}\|\leq\|A\|_{\mathrm{e}}\cdot\textit{tol},\qquad i=1,\dots,j$
for a user-prescribed stopping tolerance $\textit{tol}>0$, where $\|A\|_{\rm e}$ is defined as in (4.39). Denote
@@
[…]
@@
If $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ does not converge, we take $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=1,i_{2},\dots,i_{\tilde{m}}$ that meet (4.39) to form $\widetilde{U}_{\tilde{m}}$ and $\widetilde{V}_{\tilde{m}}$ in (4.40), as described in section 4.3. Then we set $U_{\mathrm{p}}=[U_{\mathrm{c}},\widetilde{U}_{\tilde{m}}]$ and $V_{\mathrm{p}}=[V_{\mathrm{c}},\widetilde{V}_{\tilde{m}}]$, and use MINRES to approximately solve the new preconditioned correction equation
(5.6) $\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}\!\!&\\ &\!\!I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}\!\!&\\ &\!\!I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=r_{\mathrm{p}}$
for $(s,t)\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$, where
$r_{\mathrm{p}}:=-\begin{bmatrix}I-U_{\mathrm{c}}U_{\mathrm{c}}^{T}&\\ &I-V_{\mathrm{c}}V_{\mathrm{c}}^{T}\end{bmatrix}r$
@@
with $r=r(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ the residual of $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ defined by (2.4). We compute an approximate solution $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$ and terminate MINRES when the inner residual $r_{\mathrm{in}}$ in (4.13) satisfies
(5.7) $\|r_{\mathrm{in}}\|\leq\|r\|\cdot\min\{{\rho\tilde{\varepsilon}},0.01\},$
@@
where $\tilde{\varepsilon}\in[10^{-4},10^{-3}]$ is a user-prescribed tolerance and $\rho$ is a constant that is estimated by using all the approximate singular values computed at the current cycle; see [10] for details. We use the approximate solution $\tilde{s}$ and $\tilde{t}$ to expand the subspaces $\mathcal{U}$ and $\mathcal{V}$. Since $(\widetilde{U},\widetilde{V})\perp\perp(U_{\mathrm{c}},V_{\mathrm{c}})$ and $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{c}},V_{\mathrm{c}})$, the expanded $\mathcal{U}$ and $\mathcal{V}$ are automatically orthogonal to $\mathcal{R}(U_{\mathrm{c}})$ and $\mathcal{R}(V_{\mathrm{c}})$, respectively. IPJDSVD then proceeds in a regular way until $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ converges to $(\sigma_{j+1},u_{j+1},v_{j+1})$. We denote $\left(\theta_{j+1,\mathrm{c}},u_{j+1,\mathrm{c}},v_{j+1,\mathrm{c}}\right)=(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$, add it to the converged $\left(\Sigma_{\mathrm{c}},U_{\mathrm{c}},V_{\mathrm{c}}\right)$, and set $j=j+1$. Proceed in such a way until all the $\ell$ singular triplets are found.
@@
In the deflated IPJDSVD method described above, we now have $1\leq\tilde{m}\leq m-j$ for $j<m$ since we have already computed $j$ converged singular triplets and the deflated matrix has only $m-j$ clustered singular values closest to $\tau$. For $j\geq m$, the preconditioned correction equations degenerate to the standard ones, and non-restarted IPJDSVD reduces to the standard JDSVD.
@@
For the purgation issue, we may benefit much from making full use of the current $\mathcal{U}$ and $\mathcal{V}$ when starting to compute the next singular triplet $(\sigma_{j+1},u_{j+1},v_{j+1})$, and, compared with the computation of $(\sigma_{1},u_{1},v_{1})$, save outer iterations considerably when computing $(\sigma_{i},u_{i},v_{i}),\ i=2,\ldots,\ell$. The purgation approach [10] can be straightforwardly introduced into the thick-restart IPJDSVD.
@@
[…]
@@
11:     Determine the $\tilde{m}-1$ indices $i=i_{2},\dots,i_{\tilde{m}}$ according to the condition in (4.39) and form $\widetilde{U}_{\tilde{m}}=[\tilde{u}_{1},\tilde{u}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]$ and $\widetilde{V}_{\tilde{m}}=[\tilde{v}_{1},\tilde{v}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]$.
@@
for the approximate solution $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$ with the inner residual $r_{\mathrm{in}}$ satisfying the stopping criterion (5.7).
13:     if $k=k_{\max}$ then perform the thick-restart and reset $k=k_{\mathrm{new}}$ in (5.1). fi
@@ 5.3 Thick-restart IPJDSVD algorithm with deflation and purgation
(5.8) $\sqrt{\|AV_{c}-U_{c}\Sigma_{c}\|_{\mathrm{F}}^{2}+\|A^{T}U_{c}-V_{c}\Sigma_{c}\|_{\mathrm{F}}^{2}}\leq\sqrt{\ell}\|A\|_{\mathrm{e}}\cdot\textit{tol},$
where $\|\cdot\|_{\rm F}$ is the F-norm of a matrix. Other optional parameters are normalized starting vectors $u_{0}\in\mathbb{R}^{M}$ and $v_{0}\in\mathbb{R}^{N}$, the maximum and minimum dimensions $k_{\max}$ and $k_{\min}$ of the searching subspaces, the accuracy requirement $\tilde{\varepsilon}$ for expansion vectors in (5.7), and the preconditioning parameters $\textit{pretol}_{1}$ and $\textit{pretol}_{2}$ in (4.39). By defaults, we set $tol=10^{-8}$, $u_{0}=\frac{1}{\sqrt{M}}[1,\dots,1]$, $v_{0}=\frac{1}{\sqrt{N}}[1,\dots,1]$, $k_{\max}=30$, $k_{\min}=3$, $\tilde{\varepsilon}=10^{-4}$, $\textit{pretol}_{1}=0.05$ and $\textit{pretol}_{2}=0.01$.
@@
(5.9) $\widetilde{U}=[\tilde{u}_{2},\dots,\tilde{u}_{k}],\qquad\widetilde{V}=[\tilde{v}_{2},\dots,\tilde{v}_{k}],\qquad H=\mathrm{diag}\{\theta_{2},\dots,\theta_{k}\}.$
@@ 6 Numerical experiments
We report numerical experiments to illustrate the considerable superiority of IPJDSVD to JDSVD. All the experiments were performed on a 13th Gen Intel(R) Core(TM) i9-13900KF CPU 3.00 GHz with 64 GB main memory and 24 cores using the Matlab R2022b under the Microsoft Windows 11 Pro 64-bit system with the machine precision $\epsilon=2.22\times 10^{-16}$.
For both algorithms we took the default parameters as described in section 5.3, except for JDSVD, where $\textit{pretol}_{1}=0$ and $\textit{pretol}_{2}=0$. For (2.7), (4.10) and (5.6), we took zero vector as the initial guesses and used the Matlab function minres to solve them until the relative residual norms satisfied (5.7).
In all the tables, we denote by $I_{\mathrm{out}}$ and $I_{\mathrm{in}}$ the total outer and inner iterations used by each algorithm to achieve the convergence, and by $T_{\mathrm{cpu}}$ the total CPU time in seconds counted by the Matlab commands tic and toc. Notice that solving correction equations dominates the overall efficiency of IPJDSVD and JDSVD and their cost at each outer iteration. Here $I_{\mathrm{out}}$ denotes the total number of correction equations solved, which equals the actual outer iterations minus $\ell$ because one does not solve correction equations in the final outer iteration when computing each singular triplet.
@@
degme 185501 659415 8127528   2.24e+3 4.13

### v2

@@ 5 A thick-restart IPJDSVD algorithm with deflation and purgation
For practical purpose, we must limit the subspace dimension $k\leq k_{\max}$. If JDSVD and IPJDSVD do not yet converge until $k=k_{\max}$, we must restart them. Furthermore, if we are required to compute $\ell>1$ singular triplets of $A$ corresponding to the $\ell$ singular values $\sigma_{1},\ldots,\sigma_{\ell}$ closest to the target $\tau$ (cf. (2)), an appropriate deflation is necessary. Besides, it appears that a certain purgation is crucial and may greatly reduce the outer iterations for computing the second to the $\ell$th singular triplets.
@@ 5.1 A new thick-restart
If IPJDSVD does not yet converge for computing the desired $\ell$ singular triplets for a given $k_{\max}$, we modify a commonly used thick restart, initially proposed in [24] and later popularized in the JDSVD and JDGSVD type methods [9, 10, 11, 28], and propose a more effective thick-restart IPJDSVD than the standard thick-restart JDSVD. Our thick-restart retains some $k_{\operatorname{new}}$-dimensional restarting subspaces with
(87) $k_{\operatorname{new}}=\max\{k_{\min},\tilde{m}\}$
where $k_{\min}>1$ is a user-prescribed minimal dimension of the searching subspaces and $\tilde{m}$ is the number of approximate singular triplets of $A$ that are selected to participate in the preconditioned correction equation (56) in the current expansion phase.
Specifically, assume that $\tilde{m}$ approximate singular triplets $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i}),i=1,i_{2},\dots$, $i_{\tilde{m}}$ of $A$ with $i_{2},\dots,i_{\tilde{m}}\in\{2,\dots,k\}$ join the preconditioned (56) before restart. If $\tilde{m}\leq k_{\min}$, we simply restore the standard thick-restart scheme, which forms the orthonormal basis matrices of the new $k_{\min}$-dimensional subspaces $\mathcal{U}$ and $\mathcal{V}$ by
(88) $\widetilde{U}=[\tilde{u}_{1},\tilde{u}_{2},\dots,\tilde{u}_{k_{\min}}],\qquad\widetilde{V}=[\tilde{v}_{1},\tilde{v}_{2},\dots,\tilde{v}_{k_{\min}}],$
@@
(89) $H=\widetilde{U}^{T}A\widetilde{V}=\operatorname{diag}\{\theta_{1},\theta_{2},\dots,\theta_{k_{\min}}\}.$
If $\tilde{m}>k_{\min}$, we form the orthonormal basis matrices of the new $\tilde{m}$-dimensional $\mathcal{U}$ and $\mathcal{V}$ and the corresponding projection matrix of $A$ by
(90) $\widetilde{U}=\widetilde{U}_{\tilde{m}},\qquad\widetilde{V}=\widetilde{V}_{\tilde{m}},\qquad H=\widetilde{U}_{\tilde{m}}^{T}A\widetilde{V}_{\tilde{m}}=\operatorname{diag}\{\theta_{1},\theta_{i_{2}},\dots,\theta_{i_{\tilde{m}}}\}$
with $\widetilde{U}_{\tilde{m}}$ and $\widetilde{V}_{\tilde{m}}$ defined by (86). In this way, the $k_{\rm new}$-dimensional restarting subspaces contain approximate left and right singular subspaces associated with all available approximate singular values clustered at $\tau$. Whenever $m<k_{\max}$ considerably, the new thick-restart IPJDSVD is expected to substantially outperform the standard thick-restart JDSVD in terms of the outer iterations.
@@ 5.2 Deflation and purgation
Suppose that we are required to compute the $\ell>1$ singular triplets of $A$ corresponding to the $\ell$ singular values closest to the target $\tau$. Then IPJDSVD combined with the deflation advocated in [9] meets this demand perfectly. Concretely, assume that $j<\ell$ approximate singular triplets $\left(\theta_{i,\mathrm{c}},u_{i,\mathrm{c}},v_{i,\mathrm{c}}\right)$ have already converged to $(\sigma_{i},u_{i},v_{i})$, $i=1,\dots,j$ with the residual $r^{(i,\mathrm{c})}=r(\theta_{i,\mathrm{c}},u_{i,\mathrm{c}},v_{i,\mathrm{c}})$ defined by (6) satisfying
(91) $\|r^{(i,\mathrm{c})}\|\leq\|A\|_{\mathrm{e}}\cdot\textit{tol},\qquad i=1,\dots,j$
for a user-prescribed tolerance $\textit{tol}>0$, where $\|A\|_{\rm e}$ is defined as in (85). Denote
@@
Then $\left(\Sigma_{\mathrm{c}},U_{\mathrm{c}},V_{\mathrm{c}}\right)$ is a converged approximation to the partial SVD $(\Sigma_{j},U_{j},V_{j})$ of $A$ in (48). Suppose that the next singular triplet $(\sigma_{j+1},u_{j+1},v_{j+1})$ of $A$ is desired. At the extraction phase, the deflation in [9] requires that IPJDSVD implements a chosen extraction approach, e.g., the standard extraction, with respect to the $k$-dimensional $\mathcal{U}$ and $\mathcal{V}$ that are orthogonal to $\mathcal{R}(U_{\mathrm{c}})$ and $\mathcal{R}(V_{\mathrm{c}})$, respectively; we obtain approximate singular triplets $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=1,\dots,k$ as described in Section 2, and take $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ as the approximation to $(\sigma_{j+1},u_{j+1},v_{j+1})$.
@@
If $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ does not converge, we take $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=1,i_{2},\dots,i_{\tilde{m}}$ that meet (85) to form $\widetilde{U}_{\tilde{m}}$ and $\widetilde{V}_{\tilde{m}}$ in (86). Then we set $U_{\mathrm{p}}=[U_{\mathrm{c}},\widetilde{U}_{\tilde{m}}]$ and $V_{\mathrm{p}}=[V_{\mathrm{c}},\widetilde{V}_{\tilde{m}}]$, and use MINRES to approximately solve the new preconditioned correction equation
(92) $\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}\!\!&\\ &\!\!I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}\!\!&\\ &\!\!I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=r_{\mathrm{p}}$
@@
for $(s,t)\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$, where $r_{\mathrm{p}}=-\begin{bmatrix}I-U_{\mathrm{c}}U_{\mathrm{c}}^{T}\!\!\!&\\ &\!\!\!I-V_{\mathrm{c}}V_{\mathrm{c}}^{T}\end{bmatrix}r$ with $r=r(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ the residual of $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ defined by (6). We compute an approximate solution $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$ and terminate MINRES when the inner residual $r_{\mathrm{in}}$ in (59) satisfies
(93) $\|r_{\mathrm{in}}\|\leq\|r\|\cdot\min\{{\rho\tilde{\varepsilon}},0.1\},$
@@
where $\tilde{\varepsilon}\in[10^{-4},10^{-3}]$ is a user-prescribed tolerance and $\rho$ is a constant that is estimated by using all the approximate singular values computed at the current cycle; see [9] for details. We use the approximate solution $\tilde{s}$ and $\tilde{t}$ to expand the subspaces $\mathcal{U}$ and $\mathcal{V}$. Since $(\widetilde{U},\widetilde{V})\perp\perp(U_{\mathrm{c}},V_{\mathrm{c}})$ and $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{c}},V_{\mathrm{c}})$, the expanded $\mathcal{U}$ and $\mathcal{V}$ are automatically orthogonal to $\mathcal{R}(U_{\mathrm{c}})$ and $\mathcal{R}(V_{\mathrm{c}})$, respectively. IPJDSVD then proceeds in a regular way until convergence. We then assign $\left(\theta_{j+1,\mathrm{c}},u_{j+1,\mathrm{c}},v_{j+1,\mathrm{c}}\right):=(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$, add it to the converged $\left(\Sigma_{\mathrm{c}},U_{\mathrm{c}},V_{\mathrm{c}}\right)$, and set $j=j+1$. Proceed in such a way until all the $\ell$ singular triplets are found.
@@
In the deflated IPJDSVD method described above, we now have $1\leq\tilde{m}\leq m-j$ for $j<m$ since we have already computed $j$ converged singular triplets and the deflated matrix has only $m-j$ clustered singular values near $\tau$. For $j\geq m$, the preconditioned correction equations degenerate to the standard ones, and IPJDSVD reduces to JDSVD.
@@
For the purgation issue, we may benefit much from making full use of the current $\mathcal{U}$ and $\mathcal{V}$ when starting to compute the next singular triplet $(\sigma_{j+1},u_{j+1},v_{j+1})$, and, compared to the computation of $(\sigma_{1},u_{1},v_{1})$, save outer iterations considerably when computing $(\sigma_{i},u_{i},v_{i}),\ i=2,\ldots,\ell$. The purgation approach [9] can be straightforwardly introduced into the thick-restart IPJDSVD.
@@
7:        Update $\Sigma_{\mathrm{c}}=\operatorname{diag}\{\Sigma_{\mathrm{c}},\theta_{1}\},U_{\mathrm{c}}=[U_{\mathrm{c}},\tilde{u}_{1}]$, $V_{\mathrm{c}}=[V_{\mathrm{c}},\tilde{v}_{1}]$, and set $k_{c}=k_{c}+1$.
@@
11:     Determine the $\tilde{m}-1$ indices $i=i_{2},\dots,i_{\tilde{m}}$ according to the condition in (85), and form $\widetilde{U}_{\tilde{m}}=[\tilde{u}_{1},\tilde{u}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]$ and $\widetilde{V}_{\tilde{m}}=[\tilde{v}_{1},\tilde{v}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]$.
@@
for the approximate solution $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$ with the inner residual norm $\|r_{\mathrm{in}}\|$ satisfying the stopping criterion (93).
13:     if $k=k_{\max}$ then perform the thick-restart and reset $k=k_{\operatorname{new}}$ in (87). fi
@@ 5.3 Thick-restart IPJDSVD algorithm with deflation and purgation
(94) $\sqrt{\|AV_{c}-U_{c}\Sigma_{c}\|_{\mathrm{F}}^{2}+\|A^{T}U_{c}-V_{c}\Sigma_{c}\|_{\mathrm{F}}^{2}}\leq\sqrt{\ell}\|A\|_{\mathrm{e}}\cdot\textit{tol},$
where $\|\cdot\|_{\rm F}$ is the F-norm of a matrix. Other optional parameters are normalized starting vectors $u_{0}\in\mathbb{R}^{M}$ and $v_{0}\in\mathbb{R}^{N}$, the maximum and minimum dimensions $k_{\max}$ and $k_{\min}$ of the searching subspaces, the accuracy requirement $\tilde{\varepsilon}$ for expansion vectors in (93), and the preconditioning parameters $\textit{pretol}_{1}$ and $\textit{pretol}_{2}$ in (85). By defaults, we set $tol=10^{-8}$, $u_{0}=\frac{1}{\sqrt{M}}[1,\dots,1]$, $v_{0}=\frac{1}{\sqrt{N}}[1,\dots,1]$, $k_{\max}=30$, $k_{\min}=3$, $\tilde{\varepsilon}=10^{-4}$, $\textit{pretol}_{1}=0.05$ and $\textit{pretol}_{2}=0.01$.
@@
(95) $\widetilde{U}=[\tilde{u}_{2},\dots,\tilde{u}_{k}],\qquad\widetilde{V}=[\tilde{v}_{2},\dots,\tilde{v}_{k}],\qquad H=\mathrm{diag}\{\theta_{2},\dots,\theta_{k}\}.$
@@ 6 Numerical experiments
We report numerical experiments to illustrate the considerable superiority of IPJDSVD to JDSVD. We have also developed a hybrid two-stage IPJDSVD algorithm in Matlab, which is based on the same motivation and is along the same line as that of the state-of-the-art two-stage software package PRIMME_SVDS in C language developed in [27] for accurately computing several extreme singular triplets. In the first stage of IPJDSVD_HYBRID we perform a similar IPJDSVD method on the eigenproblem of the cross product matrix $A^{T}A$ or that of $AA^{T}$ if $m<n$ with the same outer stopping tolerance as that of PRIMME_SVDS, where $A$ has full column or row rank; a convergence analysis of IPJDSVD on such eigenproblem can be made similarly and is adaptable from the current context, and we omit it due to space. We will illustrate that IPJDSVD_HYBRID is at least competitive with and can outperform PRIMME_SVDS considerably. All the experiments were implemented on a 12th Gen Intel(R) Core(TM) i7-12700K CPU 3.60GHz with 32 GB main memory and 12 cores using the Matlab R2023b under the Ubuntu 22.04.4 LTS 64 bit system with the machine precision $\epsilon=2.22\times 10^{-16}$.
@@
3elt 4720 4720 27444 6.03 2.12e-3
plddb 3069 5049 10839 1.43e+2 1.16e-2
bas1lp 5411 9825 587775 1.49e+3 3.62e-2
G66 9000 9000 36000 3.58 2.14e-4
stat96v4 3173 63076 491336 1.96e+1 2.07e-3
atmosmodl 1489752 1489752 10319760 6.21e+5 5.42e+2
cage15 5154859 5154859 99199551 1.02 8.56e-2
Table 1 lists the test matrices from the SuiteSparse Matrix Collection [2] together with some of their basic properties, where nnz denotes the total number of nonzero entries in $A$, the largest and smallest singular values $\sigma_{\max}$ and $\sigma_{\min}$ of $A$ are, only for experimental purpose, computed by the Matlab functions svds and svd for the very large matrices “relat8”, “tomographic1”, “atmosmodl”, “cage15” and all the other moderately large matrices, respectively.
In all the tables, denote by $I_{\mathrm{out}}$ and MVs the total number of correction equations and matrix-vector multiplications with $A$ and $A^{T}$ that the underlying algorithm was run for the same outer convergence tolerance, and by $T_{\mathrm{cpu}}$ the total CPU time in seconds. Notice that MVs dominates the overall efficiency of each algorithm. Since our codes were written in Matlab and the advanced PRIMME_SVDS was optimally programmed in C language, the latter should normally be much more efficient than ours in terms of CPU time if the same MVs are used.
6.1 A comparison of IPJDSVD and JDSVD
We took the default parameters as described in Section 5.3. For (9), (56) and (92), we took zero vector as the initial guesses and used the Matlab function minres to solve them until the relative residual norms met (93). To illustrate the wide applicability of our algorithms, we took $\tau$ at will, so that they correspond to both extreme and interior singular triplets.